OpenStudy (anonymous):

Vectors a, b. |a-b| I'm not sure what the first step is. Am I finding the difference in the vectors?

OpenStudy (anonymous):

If you have the vectors numerically, yes subtract b from a. Then the magnitude of the vector (the two lines) is given by the square root of the sum of the squares of each component. So if a-b came out as (3, 4, 12) , you'd get \[|a-b| = \sqrt{3^2+4^2+12^2} = \sqrt{169}=13\] Hope that makes sense!

OpenStudy (anonymous):

I think that does make sense... Let me try it. Thank you for such a detailed response!!

OpenStudy (anonymous):

Vectors a= <5, -12>, b= <-3,-6> The answer should be 10, according to the book. Could you walk me through solving the problem?

OpenStudy (anonymous):

Nevermind~ I just went over a third time and found a mathematical error. The computation makes sense. Thanks again!

OpenStudy (anonymous):

Sorry I didn't see you had responded. Glad you got it!